OpenStudy (ttop0816):

Help please! Convert the Cartesian coordinates (1, -√3) to polar coordinates. (2, 60°) (2, 300°) (4, 300°) (4, 330°)

OpenStudy (xapproachesinfinity):

you know that \(tan\theta=x/y=-1/\sqrt{3}\) now find theta

OpenStudy (xapproachesinfinity):

once you find that, plug it in x=rcostheta y=rsintheta

OpenStudy (xapproachesinfinity):

you should get your r and theta